<e> Evaluate the integral ∮c= 1 /(z-1)(z-2)(z-4) dz where C is |z| = 3 ?




1
Expert's answer
2022-02-21T15:55:55-0500

Let us evaluate the integral "∮_c\frac{dz}{(z-1)(z-2)(z-4)}" where "C" is "|z| = 3."


It follows that "z_1=1" and "z_2=2" are the simple poles containing in the inner part of the circle "|z|=3."


According to residue theorem


"∮_c\frac{dz}{(z-1)(z-2)(z-4)}=2\pi i(Res_{z=1}\frac{1}{(z-1)(z-2)(z-4)}+Res_{z=2}\frac{1}{(z-1)(z-2)(z-4)})
\\
=2\pi i(\frac{1}{(z-2)(z-4)}|_{z=1}+\frac{1}{(z-1)(z-4)}|_{z=2})
=2\pi i(\frac{1}{(1-2)(1-4)}+\frac{1}{(2-1)(2-4)})
\\=2\pi i(\frac{1}{3}-\frac{1}{2})
=2\pi i(-\frac{1}{6})=-\frac{\pi i}3."